TMS Meeting Minutes, May 7, 2021

Attendees: Andrew, Christian, Heejoo, Olga, Trenton (Zoom Meeting)

Etalon Test:
  • Fibers unplugged so that Etalon can access the mission at all times
  • As of three days ago, Etalon is still working on this
  • It seems they have not started doing laser on-off tests
  • We are currently waiting for the bino code test, so we need this to move along (Zenith mode, closed-dome testing)
  • Heejoo going to send a follow up to Etalon to encourage them to pick up the pace (would like it by next week)
Update from Yang:
  • Still working on bino code, though almost finished
  • Working on polishing the previous code with descriptive comments and better structure for GUI application
  • Developing a more intuitive, easier implementation of TMS
  • Yang proposes that he will have the TMS bino code by midway through next week
  • Assuming we have code that's working at the end of last week, Andrew wonders when will be the first time we have access to on-sky testing
  • Olga notes that there is some engineering time coming up sometime in the middle of this month
  • Engineering half night on May 20 (21st UT), but the polarimeter for PEPSI was just installed and needs to be checked on-sky
  • If flexible, there is a lot of time for commissioning SOL, probably extra, Christian doesn't mind allocating some of this time to TMS
  • There are also two nights of discretionary time and Christian willing to provide some time on LBC
  • Andrew remarks that we are looking to start a full commissioning sequence, and this would be great for us to start work on during these discretionary nights
Commissioning Plans
  • Andrew notes that we have killed the "throw focus" problem with our standard deviation filter
  • TMS failure modes are very graceful
  • See the "Commissioning plan starting point" document
  • Christian wonders how many times we have had TMS running well in active mode
  • Andrew would really like to get bino TMS up and running now
  • John remarks that we don't need to wait on the thermal model of the mirror to finish the TMS
  • Olga notes the risk is that users might expect to be able to run TMS all night and have FPIA hold focus throughout observations with no hitch
  • When we run TMS in passive mode, we obtain a database with which to compare many different observational environments
  • Able to clearly pick out TMS performance from FPIA performance with thermals, structural, etc.
  • Commissioning doesn't just involve active control but also passive mode with data processing as the primary focus
  • Andrew to circulate a document to set up a data processing pipeline for passive data so we know what we want to do with it
  • Both LSU and INAF are using LBC quite a bit, and they may have some interest in allowing us to commission TMS

Summer Break:
  • Heejoo to prepare multi-channel fiber routing for Gregorian mode operation, etc.
  • Heejoo has a document, "Fiber occupation AR version", which indicates which current LBC modes and Gregorian modes have addressible fiber channels at present
  • Heejoo suggests swapping between LBC and Gregorian mode, to obtain a full set of channel data
  • Note, this is not part of the commissioning, this is just part of future plans to expand TMS
  • All fiber is currently single-mode, no multi-mode fibers on the telescope right now
  • Andrew suggests that updating the fibers to multi-mode to use the MPO casette fiber switch

Thermal Focus Analysis:
  • Trenton has made a little more progress on this analysis
  • Interesting to note the correlation between faceplate to backplate temperature difference and LBCFPIA z4 (
-- TrentonBrendel - 07 May 2021
Topic revision: r1 - 10 May 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback